# Algorithmic Fairness amid Social Determinants: Reflection, Characterization, and Approach

### v1

@@ B.1 Background Information of Admission Strategies
Then, when the selective admission availability is limited such that $q_{o}<\frac{k_{o}\ln(\theta^{(\mathrm{poor})}/\theta^{(\mathrm{rich})})}{1/\theta^{(\mathrm{rich})}-1/\theta^{(\mathrm{poor})}}$, because of the CDF dominance of the rich region over the poor region (Assumption 3.3), and that we can equivalently compare thresholds $q_{\dagger}/\eta_{\dagger}>q_{o}>q_{\dagger}$ at the original-scale CDF $F^{(r)}(\cdot)$, we have:
@@
B.4 Proof of Theorem 3.7 in Section 3.4
Theorem (Top-Percentage Plans Reallocate Spots from Rich Region to Poor Region).
Under Assumptions 3.13.4, let us denote with $q_{o}$ the default threshold for selective admission, and with $q^{(\mathrm{poor})}$ and $q^{(\mathrm{rich})}$ the thresholds for poor and rich regions, respectively, if top-percentage plans are employed. Then, the increase in selective admissions (in terms of counts) for applicants from the poor region, comes from spots reallocated out of the rich region. This redistribution is a result of the top-percentage plans, and is not relevant to applicants’ demographic group:
@@
Compare Equations B.13 and B.14, we have:
@@
Because of the numerical property of Gamma CDF’s (as we have seen in the proof for Theorem 3.6), when region-specific shape parameters satisfy $k^{(\mathrm{poor})}=k^{(\mathrm{rich})}=k$, we have:
@@
Appendix C Additional Results and Discussions on Empirical Analyses
In this section, we present additional results and discussions on empirical experiments. In Section C.1, we provide a remark on the procedural fairness implications of different admission procedures. In Section C.2, we provide experimental details on University of California undergrad admission data, as well as further discussions of the empirical results. Then in Section D, we present additional empirical analyses based on the US Census data. The experiments are conducted on a laptop with Apple M1 Max chip and 32GB memory.
C.1 Remark on Procedural Fairness Implications of Different Admission Procedures
Although all three types of admission procedures share the goal of promoting fairness and diversity within the student body, the limited availability of selective admissions leads to varying redistributions of benefits and burdens among applicants. Quota-based admissions, while being rigid and mechanical, are more direct in reserving spots for URM applicants. However, as an unintended consequence, non-URM applicants from less well-off areas can be further disadvantaged when quota-based admissions are employed (Theorem 3.5). Holistic review with plus factors, in comparison, takes a more flexible approach when granting preferences to URM applicants. However, the increase in selective admission probability for URM applicants, which is reallocated from non-URM applicants, rewards the rich region more than the poor region (Theorem 3.6). Top-percentage plans, which provide race-neutral alternatives to the previous two affirmative-action strategies, transfer opportunities from rich region to poor region, operating in proportion to natural region-specific demographic compositions (Theorem 3.7).
The benefits and burdens experienced by applicants from different backgrounds in college admissions extend beyond whether or not and how the protected feature race is explicitly used in decision-making. Our theoretical results demonstrate the crucial role played by social determinants enclosed in Address Region for procedural fairness analysis. Without them, it is impossible to identify the newly introduced unfairness, since the address variable is absent from the causal graph in previous literature [72, 86, 28, 133].
C.2 Empirical Analyses on University of California Undergrad Admission
We provide description of the data, clarification of the Gamma parameterization for score distribution, and further discussions on the empirical results presented in Section 4.
C.2.1 Description of the Data
The University of California (UC) system is a public university system in the US. The UC Information Center provide summary statistics of undergrad admissions each year, including the undergraduate admissions summary, and the freshmen fall admissions summary. Because of legal and ethical considerations, the detailed data points at the individual level are not publicly available.
In the empirical analyses presented in Section 4, we utilize the university-wide (i.e., across the UC system) summary statistics of undergraduate admissions. Specifically, among the data for applicants (those who applied to at least one colleges in UC system), admissions (those who got offers from at least one college in UC), and enrollments (those who accepted the offers and enrolled in a specific college in UC), we utilize the application and admission statistics.
The undergraduate admissions summary provides the number of applicants and admitted students. [^8] https://www.universityofcalifornia.edu/about-us/information-center/admissions-residency-and-ethnicity For a specific year and campus, the data takes a form of breakdown-counts across different demographic groups, including African American, American Indian, Asian, Hispanic/Latino(a), Pacific Islander, White, Unknown, International. The freshmen fall admissions summary provides the proportion of applicants and admitted students whose characteristics satisfy certain conditions. [^9] https://www.universityofcalifornia.edu/about-us/information-center/freshman-admissions-summary For instance, the quantile statistics for high school weighted cumulative grade point average can be retrieved with the “HS weighted, capped GPA” option. All summary statistics are de-duplicated to avoid multiple-counting of students who applied to or admitted by multiple colleges at UC.
(a) Score densities
(b) Race- and region- specific composition
Figure 3: Recapitulation of Figures 2(a) and 2(b) in appendix, enlarged for better readability. Panel (a): region-specific and overall densities of academic preparedness. Panel (b): for each group, the region-specific compositions of application and admission proportions (left four subplots); for each region, the demographic composition of applicants (right subplot).
C.2.2 Gamma Parameterization for Score Distribution
Previous literature in educational research found that the distribution of student scores is roughly bell-shaped but is often not perfectly Gaussian (see, e.g., Arthurs et al. [8]). The distribution tends to skew towards the low-score end, and the support is often bounded (e.g., falls in $S_{\mathrm{MIN}},S_{\mathrm{MAX}}]$). Therefore, we use Gamma distributions to parameterize the score distribution, and utilize the shape and scale parameters to model the skewness and long-tail behaviors of the score distribution. This is consistent to Assumption 3.3 utilized in our theoretical analyses.
C.2.3 Formulation of the Constrained Optimization Problem
Let $\mathcal{L}(\cdot)$ denote the loss function:
[…]
Here, $q^{*}$’s are certain quantiles specified in the publicly-available statistics provided by University of California undergrad admissions summary, that (before the relative log conversion) correspond to capped and weighted high-school GPA scores $\{4.0,3.7,3.3,3.0\}$. We consider $\min(q^{*},q^{(r)})$ when calculating estimated cumulative probabilities for admissions, $F^{(r)}\big{(}\min(q^{*},q^{(r)})\big{)}$, because threshold values may differ across regions as a result of the employed admission procedure. The 1996 California Proposition 209 banned the use of race, ethnicity, and gender in public university admissions. Therefore, thresholds are (potentially) region-specific but race-irrelevant, i.e., $q^{(r)}$ instead of $q_{a}^{(r)}$.
C.2.4 Further Discussions on Empirical Results
We provide further discussions on empirical results, especially Figures 2(a) and 2(b), enlarged and recapitulated in Figure 3 for better readability. Here, the regions may not correspond to real geographical locations due to the the under-constrained nature of the optimization problem (Section 4.1), and we focus on the interpretation of the results in terms of the relation among characteristics of regions, demographic groups, and academic preparedness. In Section D, we will present data analyses based on the US Census data, where more detailed geographical information is available.
Figure 3(a) presents the region-specific densities of academic preparedness of applicants, as well as the overall density if we consider all applicants. We consider the pool of applicants, instead of that of admitted or enrolled students, since the application data is not yet “selected” by the university through the admission decision-making process, and therefore, more closely represents the underlying distribution of academic preparedness. Since the mean of a variable that follows Gamma distribution $\Gamma(k,\theta)$ is $k\cdot\theta$, the average score is 3.34 (60.03=0.186*0.03=0.18 converted back to the original scale) for Region 0 (blue), 2.82 for Region 1 (orange), and 3.34 for Region 2 (green). On average, the applicants in Region 0 and Region 2 have higher scores compared to those in Region 1, indicating the relative lack of educational resource in Region 1 (which results in overall insufficient academic preparedness). While the mean score is roughly the same for Region 0 and Region 2, the density of Region 0 is more concentrated at the high-score end compared to Region 2. From the resulting thresholds for the selective admissions, we can see that the threshold for Region 0 is more competitive than that for Region 2, which is further more competitive than that for Region 1.
In order to see the race-specific compositions of admissions indicated by the color-shaded areas under region-specific curves in Figure 3(a), we present Figure 3(b). We use the height of color-coded bars to denote the proportion of applicants that reside in specific regions, and the color-shaded part to indicate the proportion of admissions. For instance, for the African American group, the majority of applicants are from Region 1 (since the orange bar is highest in the upper-left subplot of Figure 3(b), corresponding to Region 1). Although more applications come from Region 1 ($36.06\%$ among all applicants), Region 1 appears to be the area where the educational resource is most scarce, and the relative concentration of African American applicants is more pronounced compared to other groups. The fact that the overall admission rate ($53\%$) is lowest for the African American group also corroborates with the previous observation. In other words, there is a correlation between region’s ethnicity composition and the state-of-affairs of social determinants, as indicated by the academic preparedness of applicants and the admission outcomes.
Appendix D Enhancing Census Data Product Though Linking Socioeconomic Status Indices
In this section, we present additional analyses on the US Census data [22, 23, 24] to further emphasize the importance of considering the social determinants in algorithmic fairness with concrete examples.
We retrieve the public use microdata sample (PUMS) data from the US Census Bureau [25], and provide visualizations of the age structure, racial composition, and occupation distribution in different Public Use Microdata Areas (PUMAs) in California based on the 2023 US Census PUMS data. PUMA is a geographical region smaller than counties, and the PUMA region is a strict subset of the corresponding state. Each PUMA contains at least $100,000$ residents and provides reliable, detailed demographic, economic, and housing statistics at a sub-state level while also protecting the confidentiality of respondents [25]. In order to link PUMAs to indicators of (area-level) social determinants, we also retrieve the updated Area Deprivation Index (ADI) data [^10] https://www.neighborhoodatlas.medicine.wisc.edu/ [69, 68] and the Social Vulnerability Index (SVI) data [^11] https://www.atsdr.cdc.gov/place-health/php/svi/svi-data-documentation-download.html [61, 44]. We provide the implementation at the Github repository https://github.com/zeyutang/FairnessAmidSocialDeterminants.
(a) PUMAs with different ADI levels (higher ADI indicates higher area deprivation).
(b) PUMAs with different SVI levels (higher SVI indicates higher social vulnerability).
Figure 4: Histogram of annual income for African American women residing in PUMAs with different ADI levels (top) and PUMAs with different SVI levels (bottom).
D.1 Insufficiency of (Intersectional) Sensitive Attributes When Capturing Disadvantage
In Figure 4, we present the histogram of annual income for African American women residing in PUMAs with different ADI and SVI levels. As we can see, although the demographic information reflects the intersectional characteristics of individuals (race and sex), the social determinants in different regions still play a nontrivial role in shaping the income distribution. For instance, in PUMAs with higher ADI levels, the income distribution is more skewed towards lower income levels, indicating that individuals in these areas may face more significant economic challenges compared to those in PUMAs with lower ADI levels. Similar patterns can be observed for the SVI levels, where PUMAs with higher SVI levels show a more pronounced skew towards lower income levels.
In below sections, we present the age structure, racial composition, occupation distribution, and their combinations in different PUMAs, to provide direct and concrete examples of how the social determinants (e.g., those associated to the PUMA regions) relate to algorithmic fairness.
@@ D.2 Age Structure of Population in PUMAs
Figure 6: Racial composition in various PUMA regions in California based on US Census data.
D.3 Racial Composition in PUMAs
In Figure 6, we present racial compositions across PUMAs. In the context of US Census data, “Hispanic or Latino(a)” origin is considered an ethnicity, not a race. Individuals of Hispanic or Latino(a) origin can be of any race and are often asked to identify both their race and their ethnicity during the data collection. Therefore, the racial composition does not contain a separate category for Hispanic or Latino(a) individuals.
@@ D.3 Racial Composition in PUMAs
Figure 7: Occupational structure in various PUMA regions in California based on US Census data.
D.4 Occupation Distribution in PUMAs
@@
D.5 Combination of Factors in PUMA
In Figure 8, we present how PUMAs can have very different profiles in terms of residents’ age structure, race decomposition, and occupation distribution. In terms of the age structure, PUMAs 3749 and 8504 show more concentrations in the 20–40 age range, while PUMA 1700 has a high proportion of senior adults. In terms of the race decomposition, the majority of residents are white ($75.9\%$) for PUMA 1700, African American ($41.5\%$), and Asian ($70.7\%$) for PUMA 8504. In terms of the occupation distribution, while the proportion of medical and healthcare practitioners is similar across the three regions, the occupational structures are very different. For instance, nearly one half of the working force in PUMA 8504 is within the category of computer and mathematical occupations, while the number is significantly lower in PUMAs 1700 and 3749, with a proportion of $18.4\%$ and $5.8\%$, respectively. The comprehensive understanding of the social determinants in different regions can help inform policy-making and resource allocation decisions, so that we can achieve algorithmic fairness in a more principled and transparent way.
@@ D.5 Combination of Factors in PUMA
[…]
[supp]

### v2

@@ B.3 Additional Theoretical Result Related to Holistic Review with Plus Factors
Then, when the selective admission availability is limited such that $q_{o}<\frac{k_{o}\ln(\theta^{(\mathrm{poor})}/\theta^{(\mathrm{rich})})}{1/\theta^{(\mathrm{rich})}-1/\theta^{(\mathrm{poor})}}$, because of the CDF dominance of the rich region over the poor region (Assumption B.1), and that we can equivalently compare thresholds $q_{\dagger}/\eta_{\dagger}>q_{o}>q_{\dagger}$ at the original-scale CDF $F^{(r)}(\cdot)$, we have:
@@
Theorem B.2 characterizes different levels of benefits for URM applicants from different regions. Specifically, in terms of the increase in admission probability to the selective college, URM applicants from the rich region benefit more from the admission procedure that utilizes holistic review with plus factors, compared to URM applicants from the poor region. To better demonstrate our theoretical result, we provide illustrations in Figure 3(b).
As presented in top-row subfigures in Figure 3(b), at the original scale, the region-specific distributions of academic preparedness are the same for URM and non-URM applicants (Assumption 4.2). Holistic review with plus factors grants preference to URM applicants by perceiving their scores, at the distribution level, as if they were sampled from a distribution that is more concentrated at the high-score end (the plus-factor scale). Because of the limited availability in selective admissions, the threshold $q_{\dagger}$ for admission under holistic review with plus factors is more competitive than the default $q_{o}$, i.e., $q_{\dagger}<q_{o}$, for both URM and non-URM applicants. While non-URM applicants are assessed on the original scale, URM applicants are evaluated on a plus-factor scale. Under the Gamma parameterization (Assumption B.1), this is equivalent to employing a more competitive threshold $q_{\dagger}$ for non-URM applicants but a less competitive one $q_{\dagger}/\eta_{\dagger}$ for URM applicants, where $q_{\dagger}<q_{o}<q_{\dagger}/\eta_{\dagger}$. Although the mathematical form of $q_{o}<\nicefrac{{k_{o}\ln(\theta^{(\mathrm{poor})}/\theta^{(\mathrm{rich})})}}{{(\nicefrac{{1}}{{\theta^{(\mathrm{rich})}}}-\nicefrac{{1}}{{\theta^{(\mathrm{poor})}}})}}$ appears convoluted, the condition itself is relatively mild. Graphically speaking, the spots at the selective college are limited such that the threshold $q_{o}$ does not reach the point where region-specific Gamma density curves (in the original scale) intersect, as depicted by $\widetilde{q}$ in Figure 3(b).
From the shaded areas in bottom-row subfigures in Figure 3(b), we can see that the increased admission probability for URM groups comes with a corresponding reduction in that for non-URM groups. However, such redistribution benefits URM applicants in the rich region more than those in the poor region, essentially disadvantaging URM applicants in less socio-economically advantaged areas.
B.4 Additional Theoretical Result Related to the Top-Percentage Plans
Taking into account the demographic composition of applicants and the number of available spots at the selective college, we characterize the difference between top-percentage plans compared to the default selective admission. When explicitly considering the role of Region in applicants’ academic preparedness, we show that the redistribution of limited selective admissions, as implied by top-percentage plans, is carried out by reallocating availability from the rich region to the poor region, regardless of the demographic group of applicants:
Theorem B.3 (Top-Percentage Plans Reallocate Spots from Rich Region to Poor Region).
Under Assumptions 4.14.4 and also B.1, let us denote with $q_{o}$ the default threshold for selective admission, and with $q^{(\mathrm{poor})}$ and $q^{(\mathrm{rich})}$ the thresholds for poor and rich regions, respectively, if top-percentage plans are employed. Then, the increase in selective admissions (in terms of counts) for applicants from the poor region, comes from spots reallocated out of the rich region. This redistribution is a result of the top-percentage plans, and is not relevant to applicants’ demographic group:
$\begin{split}\big(n_{a}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{poor})}\big)\big[F^{(\mathrm{poor})}(q^{(\mathrm{poor})})-F^{(\mathrm{poor})}(q^{(o)})\big]=\big(n_{a}^{(\mathrm{rich})}+n_{a^{\prime}}^{(\mathrm{rich})}\big)\big[F^{(\mathrm{rich})}(q^{(o)})-F^{(\mathrm{rich})}(q^{(\mathrm{rich})})\big].\end{split}$
@@ B.4 Additional Theoretical Result Related to the Top-Percentage Plans
Compare Equation (B.13) and Equation (B.14), we have:
@@
Because of the numerical property of Gamma CDF’s (as we have seen in the proof for Theorem B.2), when region-specific shape parameters satisfy $k^{(\mathrm{poor})}=k^{(\mathrm{rich})}=k$, we have:
@@
Theorem B.3 characterizes the reallocation of the selective admission spots performed by top-percentage plans. In Figure 3(c), we use shaded areas to illustrate the transfer of admission opportunity (in terms of the region-wise probability of selective admission) from the rich region to the poor region. The additional selective admissions gained by the poor region, compared to default setting, are distributed proportionally to the natural demographic composition of each group.
Appendix C Additional Data Analyses on U.S. Census Data
In this section, we present the age structure, racial composition, occupation distribution in different PUMAs.
C.1 Age Structure of Population in PUMAs
@@
C.2 Racial Composition in PUMAs
Figure 5: Racial composition in various PUMA regions in California based on U.S. Census data.
In Figure 5, we present racial compositions across PUMAs. In the context of U.S. Census data, “Hispanic or Latino(a)” origin is considered an ethnicity, not a race. Individuals of Hispanic or Latino(a) origin can be of any race and are often asked to identify both their race and their ethnicity during the data collection. Therefore, the racial composition does not contain a separate category for Hispanic or Latino(a) individuals.
@@ C.2 Racial Composition in PUMAs
C.3 Occupation Distribution in PUMAs
Figure 6: Occupational structure in various PUMA regions in California based on U.S. Census data.
@@ C.3 Occupation Distribution in PUMAs
C.4 Combination of Demographic Factors in PUMA
@@
In Figure 7, we present how PUMAs can have very different profiles in terms of residents’ age structure, racial decomposition, and occupation distribution. In terms of the age structure, PUMAs 3749 and 8504 show more concentrations in the 20–40 age range, while PUMA 1700 has a high proportion of senior adults. In terms of the race decomposition, the majority of residents are white ($75.9\%$) for PUMA 1700, African American ($41.5\%$), and Asian ($70.7\%$) for PUMA 8504. In terms of the occupation distribution, while the proportion of medical and healthcare practitioners is similar across the three regions, the occupational structures are very different. For instance, nearly one half of the working force in PUMA 8504 is within the category of computer and mathematical occupations, while the number is significantly lower in PUMAs 1700 and 3749, with a proportion of $18.4\%$ and $5.8\%$, respectively. Therefore, in addition to sensitive attributes, the comprehensive understanding of the social determinants in different regions can help inform policy-making and resource allocation decisions, so that we can achieve algorithmic fairness by taking into account both sensitive attributes and social determinants.